Alt Company is in the process of analyzing their accounts in order to prepare adjusting entries for the fiscal year 2018, which ends on December 31, 2018. The following information has been collected by the accounting staff in order to prepare the adjusting entries for 2018.

  1. The balance in the Insurance Expense account contains the total premium costs of a policy with an original cost of $10,800 that has a 3-yr. term beginning on October 1, 2018. The policy was paid for in full on October 1, 2018 and the bookkeeper recorded the full premium to the Insurance Expense account.  
  2. On July 31, 2018, Alt received $21,600 rent from its lessee for an eighteen month lease beginning August 1, 2018. The bookkeeper recorded the total $21,600 to the Rent Revenue account.
  1. On December 28, 2018, the bookkeeper credited Sales Revenue for receipt of a payment from a credit customer who was paying the $20,000 balance on their account for goods that they had purchased on account earlier in the year. The entry for the sales to the credit customer were recorded correctly at the time the goods were sold.
  2. At December 31, 2018, salaries and wages earned but not paid were $4,200.
  3. On April 30, 2018, Alt leased a new office suite for their corporate executives in a downtown building for $3,000 per month, paying rent for 12 months (May 1, 2018 through April 30, 2019) in advance. The bookkeeper recorded the full amount of the payment to Prepaid Rent.
  4. The 2018 depreciation for the administrative office equipment totals $8,000 and has not been recorded.

Required Please show all calculations

Prepare the adjusting journal entries to record the corrections and adjustments required by the information above (show calculations and omit explanations)
Answer #1
Alt Company
Journal entries
Entries Date Account Debit Credit Calculation
Correction 31-Dec Prepaid Insurance $     10,800
Insurance expense $     10,800
Adjusting 31-Dec Insurance expense $           900 =10,800/36*3
Prepaid Insurance $           900
Correction 31-Dec Rent Revenue $     21,600
Unearned Rent Revenue $     21,600
Adjusting 31-Dec Unearned Rent Revenue $        6,000 =21,600/18*5
Rent Revenue $        6,000
Correction 31-Dec Sales Revenue $     20,000
Accounts Receivable $     20,000
Adjusting 31-Dec Salaries and wages expense $        4,200
Salaries and wages Payable $        4,200
Adjusting 31-Dec Rent expense $     24,000 =3,000*8
Prepaid Rent $     24,000
Adjusting 31-Dec Depreciation expense $        8,000
Accumulated Depreciation $        8,000